# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v2

@@ 6 Conclusion
In this paper, we systematically evaluate various defense methods, both theoretically and empirically. We established the algorithmic upper bound and matching information-theoretic lower bound of the reconstruction error for various defenses on two-layer random nets, thereby quantitatively analyzing the effect of defense strength, data dimension, and model width. To extend our theoretical insights to more general network architectures and explore the utility-privacy trade-off, we introduced a robust attack method capable of overcoming a broad spectrum of defense strategies in a setting with an honest but curious server. The method enhances the conventional gradient inversion attack by combining it with the powerful feature reconstruction attack that achieves our algorithmic upper bound. We propose to evaluate each attack based on their impact on utility and effectiveness against the strongest corresponding attack. Our evaluation indicates gradient pruning is the strongest among our considered defenses and against our considered attacks, which updates some previous evaluation results that are purely based on gradient inversion attacks (Huang et al., 2021). Our work establishes a more comprehensive understanding of the data reconstruction problem in federated learning. We anticipate that our proposed framework will encourage further research and additional evaluations building on our findings.
Acknowledgement
This work was supported by the U.S. Department of Energy, Office of Science, under Award Number DE-SC0024721 and the Whitehead Fellowship from New York University. QL would like to thank Yanjun Han for the valuable discussions on the lower bound formulation of the paper.
@@ D.1 The Minimax Risk
The risk defined above provides a lower bound for the minimax expected risk of all attacking algorithms. As an information-theoretical or statistical minimax risk, we model the observation with a random noise. In our following analysis, $O$ is interpreted as model gradients with a random Gaussian noise $O(S)=\nabla L(S;\Theta)+\epsilon$. The expectation in 27 is taken over the random noise $\epsilon$.
@@ Appendix F Additional experimental details for our proposed attack
The attack minimize the objective function given in Eq. 4. We search $\alpha_{\rm TV}$ in $\{0,0.001,0.005,0.01,0.05,0.1,0.5\}$ and $\alpha_{\rm BN}$ in $\{0,0.0005,0.001,0.01,0.05,0.01\}$ for attack without defense, and apply the best choices $\alpha_{\rm TV}=0.01$ and $\alpha_{\rm BN}=0.001$ for all defenses (we did not tune $\alpha_{\rm TV}$ and $\alpha_{\rm BN}$ for each defense as we observe the proposed attack method is robust in terms of these two hyper-parameters). We optimize the attack for 10,000 iterations using Adam [Kingma and Ba, 2014], with the initial learning rate set to $0.00001$. We decay the learning rate by a factor of $0.1$ at $3/8$, $5/8$, and $7/8$ of the optimization. For feature matching loss strength $\alpha_{f}$, we search in $\{0.01,0.05,0.1,0.5\}$ for each defense and find that $0.1$ is the optimal choice.

### v3

@@ 6 Conclusion
In this paper, we systematically evaluate various defense methods, both theoretically and empirically. We established the algorithmic upper bound and matching information-theoretic lower bound of the reconstruction error for various defenses on two-layer random nets, thereby quantitatively analyzing the effect of defense strength, data dimension, and model width. To extend our theoretical insights to more general network architectures and explore the utility-privacy trade-off, we introduced a robust attack method capable of overcoming a broad spectrum of defense strategies in a setting with an honest but curious server. The method enhances the conventional gradient inversion attack by combining it with the powerful feature reconstruction attack that achieves our algorithmic upper bound. We propose to evaluate each attack based on their impact on utility and effectiveness against the strongest corresponding attack. Our evaluation indicates gradient pruning is the strongest among our considered defenses and against our considered attacks, which updates some previous evaluation results purely based on gradient inversion attacks (Huang et al., 2021). Our work establishes a more comprehensive understanding of the data reconstruction problem in federated learning. We anticipate that our proposed framework will encourage further research and additional evaluations building on our findings.
Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science Energy Earthshot Initiative as part of the project “Learning reduced models under extreme data conditions for design and rapid decision-making in complex systems" under Award #DE-SC0024721.
@@ D.1 The Minimax Risk
The risk defined above provides a lower bound for the minimax expected risk of all attacking algorithms. As an information-theoretical or statistical minimax risk, we model the observation with a random noise. In our following analysis, $O$ is interpreted as model gradients with a random Gaussian noise $O(S)=\nabla L(S;\Theta)+\epsilon$. The expectation in 26 is taken over the random noise $\epsilon$.
@@ Appendix F Additional experimental details for our proposed attack
The attack minimize the objective function given in Eq. 4.2. We search $\alpha_{\rm TV}$ in $\{0,0.001,0.005,0.01,0.05,0.1,0.5\}$ and $\alpha_{\rm BN}$ in $\{0,0.0005,0.001,0.01,0.05,0.01\}$ for attack without defense, and apply the best choices $\alpha_{\rm TV}=0.01$ and $\alpha_{\rm BN}=0.001$ for all defenses (we did not tune $\alpha_{\rm TV}$ and $\alpha_{\rm BN}$ for each defense as we observe the proposed attack method is robust in terms of these two hyper-parameters). We optimize the attack for 10,000 iterations using Adam [Kingma and Ba, 2014], with the initial learning rate set to $0.00001$. We decay the learning rate by a factor of $0.1$ at $3/8$, $5/8$, and $7/8$ of the optimization. For feature matching loss strength $\alpha_{f}$, we search in $\{0.01,0.05,0.1,0.5\}$ for each defense and find that $0.1$ is the optimal choice.
@@ Appendix I Additional experimental results
Effectiveness of random initialization.
We conduct additional experiments by attacking the model at different time steps. Specifically, the RMSE for attacks after 1, 2, and 3 epochs of training are as follows: 0.15 (0.04), 0.17 (0.09), and 0.22 (0.08), respectively. These results indicate that the attack becomes progressively more challenging as the training continues. We use the attack at random initialization in all other experiments, which is the strongest setting.